# m-KAILIN: Knowledge-Driven Agentic Scientific Corpus Distillation Framework for Biomedical Large Language Models Training

### v1

@@ 2.5 Multi-Agent Collaborative Framework
Step-3 Direct Preference Optimization: The Direct Preference Optimization (DPO) objective explicitly aligns the model’s output distribution with automated preference signals. Given a set of preference data $\{(d_{i},q^{+}_{i},q^{-}_{i})\}$, where $q^{+}_{i}$ is the preferred question over $q^{-}_{i}$ for document $d_{i}$, DPO fine-tunes the general LLM $\theta^{b}$ to maximize the likelihood of generating preferred questions while minimizing the likelihood of dispreferred ones. Formally, the DPO objective is:
@@
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\beta\log P_{\theta}(q^{+}_{i}|d_{i}))}{\exp(\beta\log P_{\theta}(q^{+}_{i}|d_{i}))+\exp(\beta\log P_{\theta}(q^{-}_{i}|d_{i}))},$ (7)
@@
Step-4 Training Corpus Dataset Construction: The optimized Question Generation Agent $\theta^{*}$ produces new question-context pairs from raw documents. After that, the Answer Generation Agent $\varphi$ provides answers, resulting in two types of ideal datasets: $\mathcal{I}_{\textsc{cpt}}=\{(q_{j},c_{j})\}_{j=1}^{N}$: For continued pre-training (CPT), consisting of question-context pairs. $\mathcal{I}_{\textsc{sft}}=\{(q_{j},c_{j},a_{j})\}_{j=1}^{N}$: For supervised fine-tuning (SFT), consisting of question-context-answer triples.
This collaborative multi-agent pipeline enables scalable, high-quality question answering data construction without reliance on manual annotation.
2.5.1 Training for Downstream Tasks
@@
Stage-1 Continuous Pre-training: The target LLM is first further pre-trained on the heuristic question-context pairs $\mathcal{I}_{\textsc{cpt}}=\{(q_{j},c_{j})\}_{j=1}^{N}$ generated by the optimized question generator. This stage aims to improve the model’s foundational biomedical knowledge and enhance its ability to understand biomedical contexts and question styles. The prompt of this stage is given as Figure 2.
@@
Stage-2 Supervised Fine-tuning: Subsequently, the model is fine-tuned on question-context-answer triples $\mathcal{I}_{\textsc{sft}}=\{(q_{j},c_{j},a_{j})\}_{j=1}^{N}$, as well as optionally on benchmark QA datasets such as PubMedQA PQA-A [30] training set. As shown in Figure 3, this prompt is designed to match the downstream QA format. This stage explicitly optimizes the model for accurate answer generation and robust question understanding.
@@
Training Procedure: In practice, the continuous pre-training step adopts the standard language modeling objective, while the supervised fine-tuning step minimizes the cross-entropy loss between the model’s predicted answers and ground-truth answers. The two-stage training pipeline ensures that the target LLM can effectively leverage both the large-scale synthetic data from m-KAILIN and high-quality human-annotated benchmarks, substantially boosting downstream biomedical QA performance.
3 Results
3.1 Experimental Setups
Base Models We utilized as the base models of each agent in our primary experiment, while also incorporating BioMistral [9] as the backend of the Question Generation Agent for building the preference dataset in training the question generator. We adopted GPT-4o as the backend of the Answer Generation Agent.
Baselines. We conducted a comprehensive evaluation of various open-source models, including LLaMA-2 [31], LLaMA-3 [29], Mistral [32], GLM-4 [33], Qwen2.5 [34], and Gemma[35], Deepseek [36], as well as proprietary models like GPT-4[27] (with MedPrompt [37, 38]), and Med-PaLM [39]. In particular, we focused on models specifically trained for the biomedical domain, such as BioMistral [9], PMC-LLaMA [12], HEAL [40], KAILIN [20] (with Qwen2.5, GLM-4, and LLama-3 as backend), and MMedLM [41], to demonstrate the effectiveness.
Evaluation Datasets. We validated the results of our main experiment using the PubMedQA benchmark [30], a dataset specifically designed to assess the performance of question-answering systems in the biomedical domain. PubMedQA is tailored to address questions relevant to biomedical literature, making it highly suitable for assessing our framework’s adaptation in this field. Additionally, we categorized the benchmark based on Medical Subject Headings (MeSH) [26] and publication dates, enabling us to evaluate our system’s improved understanding and robustness across diverse biomedical topics and varying time spans:
Platform Information. Each experiment, either utilizing the API query or performed on an Ubuntu 18.04.6 LTS system featuring an AMD EPYC 7742 processor and four NVIDIA V100 GPUs, was carried out within a Python 3.11.0 environment using PyTorch 2.1.1.
3.2 Main Results and Analysis
@@
As illustrated in Figure 4, we benchmarked m-KAILIN and various baseline models across two parameter scales: models with fewer than 13B parameters and those with 70B parameters or greater. A clear positive correlation exists between model size and performance: models with over 70B parameters generally surpass smaller models in accuracy. However, our results reveal that the application of m-KAILIN narrows this gap, allowing smaller models to achieve competitive performance. Notably, m-KAILIN(LLaMA-3-8B) nearly matches or outperforms several larger domain-specific models, underscoring the potential of architectural and training innovations to offset limitations in model scale partially. The key driver is the MeSH-based knowledge hierarchy similarity evaluation integrated within Question Evaluation Agent, which substantially enhances a model’s ability to comprehend and reason about the complex terminologies and conceptual structures of biomedical texts. This demonstrates the effectiveness of our method in adapting smaller, general-purpose models to specialized domains—a critical advantage given the rapid iteration cycles in large model development. In the large-model group, m-KAILIN significantly boosts the performance of LLaMA-3-70B, enabling it to outperform powerful closed-source systems such as GPT-4 with MedPrompt and Med-PaLM-2, despite their larger parameter counts. This suggests that, while increasing model size typically yields better results, targeted preference alignment using knowledge hierarchy evaluation can enable smaller or open-source models to close the gap or even exceed the performance of much larger models, thus offering substantial efficiency and accessibility benefits. In summary, the m-KAILIN framework enables both small and large general-purpose models to achieve state-of-the-art performance on domain-specific biomedical tasks. These experimental results also validate the feasibility of leveraging the m-KAILIN framework to automatically extract high-quality training data from raw textual sources.
3.3 Scaling Law of Dataset Distillation
Figure 5: Evaluation(accuracy(%)) of Llama-3-8B and Llama-3-70B on PubMedQA using datasets of varying scales, from 20,000 (20K) to 100,000 (100K).
We conducted experiments using datasets of varying scales for continuous pre-training, followed by supervised fine-tuning under consistent experimental settings, to evaluate the impact of different data volumes on model performance. In this study, we selected the Llama-3-8B and Llama-3-70B models as base models, to verify the impact of varying dataset scales on models of different sizes.
As shown in Figure 5, we observed that the larger LLaMA-3-70B model demonstrates overall better performance compared to the smaller LLaMA-3-8B model. We attributed this advantage to its larger parameter scale, which provides superior feature capture and generalization capabilities. We also observed that on datasets of varying sizes, both models showed a trend of performance improvement as the dataset size increased. The underlying driver is that larger datasets distilled by m-KAILIN are better able to represent a broader range of biomedical knowledge due to a greater number of documents being comprehensively analyzed. This phenomenon indicates that the m-KAILIN framework indeed produces higher-quality datasets when distilling larger-scale datasets. Moreover, we found that while smaller models may not outperform larger models, they seem to benefit more from progressively larger distilled datasets. We attribute this phenomenon to the degree of alignment between task complexity and model scale. This phenomenon indicates that larger models, due to potential knowledge redundancy, may gain less from incremental datasets on a specific task.
3.4 Component Study
@@ 3.4.1 Component Study on Question Generation Agent
3.4.2 Component Study on Context Retrieval Agent and Question Evaluation Agent
@@
Figure 7 shows the performance across different configurations. Removing MeSH (w/o MeSH) leads to a substantial decrease in accuracy, especially in the reasoning-required setting. This suggests that biomedical hierarchical knowledge—provided by MeSH—is crucial for the evaluation agent to accurately judge the relevance and quality of context-question pairs. Without this structured guidance, the general LLM struggles to evaluate biomedical content effectively, resulting in lower overall performance.
In the w/o Embedding setting, where the retrieval agent uses general-purpose BERT embeddings instead of a fine-tuned, domain-specific model, performance also drops. This indicates that domain-adapted embeddings significantly enhance the retrieval of relevant biomedical contexts. General embeddings may miss domain-specific nuances, leading to suboptimal context selection and thereby affecting downstream QA accuracy.
Notably, when both MeSH and domain-adapted embedding are removed (w/o Both), there is a pronounced performance degradation (64.8% in reasoning-required, 43% in question-only), confirming that both hierarchical knowledge in evaluation and domain adaptation in retrieval play complementary, essential roles.
Overall, these results demonstrate that integrating domain-specific components—both in evaluation and retrieval—is vital for constructing a high-quality biomedical corpus and maximizing LLM downstream performance.
3.5 Case Study of Generated Corpus
@@
From the specific cases illustrated in Figure 8, it is evident that the model can recognize ”activity” in relation to papain’s function as an enzyme, demonstrating a general understanding facilitated by domain cues. However, without comprehensive and structured biomedical knowledge, the model still exhibits limitations in generating integrative questions that require a deeper or more holistic perspective.
Performance differences under the question-only inference setting further highlight these shortcomings. In particular, when the embedding model is not fine-tuned on biomedical data (i.e., using a general-purpose embedding), the model struggles to capture question-specific nuances, resulting in significantly reduced performance when relevant context is unavailable.
Our case analysis also reveals the pivotal role of the MeSH-based knowledge hierarchy. When MeSH guidance is present, the model is able to correctly interpret papain as a ”proteolytic enzyme” within the broader biomedical context. However, the absence of fine-grained, domain-specific knowledge—especially when leveraging only general LLMs or embeddings—leads to missed opportunities in question generation, such as failing to inquire about detailed aspects like enzyme activity or dimerization.
These findings underscore the necessity of integrating both structured biomedical ontologies (e.g., MeSH) and domain-adapted retrieval mechanisms in our multi-agent framework. Only through this combination can the model attain both the breadth and depth required for high-quality, knowledge-rich biomedical question generation and evaluation.
3.6 Robustness Checking
3.6.1 Chronological Robustness.
Time Span Prop
1989-2000 9.6%
2001-2004 12.2%
2005-2007 11.9%
2008-2009 11.9%
2010-2011 9.6%
2012-2013 14.8%
2014-2015 15.0%
2016-2017 9.2%
Figure 9: Evaluations (accuracy (%)) for different chronological subsets problems in PubMedQA PQA-L on our models compared to other representative models.
As shown in Figure 9, our multi-agent framework, especially when combined with MeSH-based evaluation, enables models to maintain consistently high accuracy across documents published at different time periods. Compared to strong biomedical baselines such as MMedLM, models distilled with m-KAILIN demonstrate significantly greater robustness to temporal shifts. This improvement stems from two key innovations: (1) the multi-agent setup, which incorporates diverse perspectives during corpus construction, and (2) the MeSH-guided evaluation agent, which provides structured biomedical context, allowing the model to better understand and adapt to evolving domain knowledge over time. Notably, larger models exhibit even greater temporal robustness, likely due to their increased capacity to absorb and generalize from temporally diverse data curated by our framework.
3.6.2 Subdisciplinary Robustness.
MeSH term Samples
Female 785
Male 703
Middle Aged 542
Aged 414
Adult 492
Adolescent 204
Similarly, our evaluation on subdisciplinary slices, based on major MeSH terms (Figure 10), shows that m-KAILIN-trained models outperform their peers across various biomedical topics and demographics. This robustness arises from our MeSH-based multi-agent design: by leveraging MeSH hierarchies during both retrieval and evaluation, the framework ensures comprehensive coverage and nuanced understanding across biomedical subfields. The evaluation agent’s use of MeSH guides more relevant and context-aware question-answer generation and systematically exposes the model to the breadth of biomedical literature structure. Some variability across demographic terms highlights ongoing challenges. Still, the overall trend affirms the effectiveness of our approach in producing models well-adapted to the complex, hierarchical landscape of biomedical knowledge.
@@ 3.7 Hyperparameter Experiment
4 Related Work
4.1 Dataset Distillation.
4.2 Synthesis Text Data in Biomedical field.
4.3 LLMs in Biomedical Domain.
Recent advancements in biomedical language models have focused on improving domain adaptation to better handle dense terminology and complex medical concepts. Approaches such as retrieval-augmented generation have enhanced models’ ability to retrieve relevant information for domain-specific tasks[55, 56, 57]. PMC-LLaMA[12] has been fine-tuned using millions of biomedical papers and medical textbooks, integrating medical knowledge graphs like UMLS for domain knowledge. Similarly, BioMistral[9] has applied extensive data processing and adaptive pretraining on PubMed Central, while HEAL[40] has combined public and proprietary datasets for continuous pretraining on a general-purpose model, further enhancing its biomedical capabilities.

### v2

@@ 3.5.4 Direct Preference Optimization
The Direct Preference Optimization (DPO) objective explicitly aligns the model’s output distribution with automated preference signals. Given a set of preference data $\{(d_{i},q^{+}_{i},q^{-}_{i})\}$, where $q^{+}_{i}$ is the preferred question over $q^{-}_{i}$ for document $d_{i}$, DPO fine-tunes the general LLM $\theta^{b}$ to maximize the likelihood of generating preferred questions while minimizing the likelihood of dispreferred ones. Formally, the DPO objective is:
@@
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\left(\beta\log P_{\theta}(q_{i}^{+}|d_{i})\right)}{\begin{aligned} &\exp\left(\beta\log P_{\theta}(q_{i}^{+}|d_{i})\right)\\ &+\exp\left(\beta\log P_{\theta}(q_{i}^{-}|d_{i})\right)\end{aligned}}$ (8)
@@
Figure 4: The fine-tuned Question Generation Agent will collaorated with Answer Generation Agent and Context Retrieval Agent to build the training corpora dataset.
3.5.5 Training Corpus Dataset Construction
As shown in Figure 4, the optimized Question Generation Agent $\theta^{*}$ produces new question-context pairs from raw documents. After that, the Answer Generation Agent $\varphi$ provides answers, resulting in two types of ideal datasets: $\mathcal{I}_{\textsc{cpt}}=\{(q_{j}^{*},\textbf{c}_{j})\}_{j=1}^{N}$, for continued pre-training (CPT), consisting of question-context pairs. $\mathcal{I}_{\textsc{sft}}=\{(q^{*}_{j},\textbf{c}_{j},a^{*}_{j})\}_{j=1}^{N}$, for supervised fine-tuning (SFT), consisting of question-context-answer triples. This collaborative multi-agent pipeline enables scalable, high-quality question answering data construction without reliance on manual annotation.
3.6 Training for Downstream Tasks
@@
{bclogo}
[…]
Motivated by this study, I conducted a literature review to gather additional resources and contextualize its findings. During this process, I identified the following key materials: {Retrived Context}.
Reflecting on these insights, I formulated the following research question: {Question}.
@@
{bclogo}
[…]
Title: {Title}
Context: {Context}
Response: Question
@@
3.6.1 Continuous Pre-training
The target LLM is first further pre-trained on the heuristic question-context pairs $\mathcal{I}_{\textsc{cpt}}=\{(q_{j},\textbf{c}_{j})\}_{j=1}^{N}$ generated by the optimized question generator. This stage aims to improve the model’s foundational biomedical knowledge and enhance its ability to understand biomedical contexts and question styles. The prompt of this stage is given as Figure 5.
@@ 3.6.1 Continuous Pre-training
3.6.2 Supervised Fine-tuning
Subsequently, the model is fine-tuned on question-context-answer triples $\mathcal{I}_{\textsc{sft}}=\{(q_{j},\textbf{c}_{j},a_{j})\}_{j=1}^{N}$, as well as optionally on benchmark QA datasets such as PubMedQA PQA-A [40] training set. As shown in Figure 6, this prompt is designed to match the downstream QA format. This stage explicitly optimizes the model for accurate answer generation and robust question understanding.
3.6.3 Summary of the Training Procedure
The two-stage training pipeline ensures that the trained LLM can effectively leverage both the large-scale synthetic data from m-KAILIN and high-quality human-annotated benchmarks, substantially boosting downstream biomedical QA performance. In practice, the continuous pre-training step adopts the standard language modeling objective, while the supervised fine-tuning step minimizes the cross-entropy loss between the model’s predicted answers and ground-truth answers.
4 Results
4.1 Experimental Setups
4.1.1 Base Models
We utilized the llama-series LLM as base models in our primary experiment, while also incorporating BioMistral [14] as the backend of the Question Generation Agent for building the preference dataset and fine-tuning the question generator. We adopted GPT-4o as the backend of the Answer Generation Agent.
4.1.2 Baselines
We conducted a comprehensive evaluation of various open-source models, including LLaMA-2 [57], LLaMA-3 [56], Mistral [58], GLM-4 [59], Qwen2.5 [60], and Gemma[61], Deepseek [62], as well as proprietary models like GPT-4[54] (with MedPrompt [63, 64]), and Med-PaLM [65]. In particular, we focused on models specifically trained for the biomedical domain, such as BioMistral [14], PMC-LLaMA [17], HEAL [48], KAILIN [26] (with Qwen2.5, GLM-4, and LLama-3 as backend), and MMedLM [66], to demonstrate the effectiveness.
4.1.3 Evaluation Datasets
We validated the results of our main experiment using the PubMedQA benchmark [40], a dataset specifically designed to assess the performance of question-answering systems in the biomedical domain. PubMedQA is tailored to address questions relevant to biomedical literature, making it highly suitable for assessing our framework’s adaptation in this field. Additionally, we categorized the benchmark based on Medical Subject Headings (MeSH) [52] and publication dates, enabling us to evaluate our system’s improved understanding and robustness across diverse biomedical topics and varying time spans:
4.1.4 Platform Information
Each experiment, either utilizing the API query or performed on an Ubuntu 18.04.6 LTS system featuring an AMD EPYC 7742 processor and four NVIDIA V100 GPUs, was carried out within a Python 3.11.0 environment using PyTorch 2.1.1.
4.2 Main Results and Analysis
@@
As illustrated in Figure 7, we benchmarked m-KAILIN and various baseline models across two parameter scales: models with fewer than 13B parameters and those with 70B parameters or greater. A clear positive correlation exists between model size and performance: models with over 70B parameters generally surpass smaller models in accuracy. However, our results reveal that the application of m-KAILIN narrows this gap, allowing smaller models to achieve competitive performance. Notably, m-KAILIN-LLaMA-3 (8B) nearly matches or outperforms several larger domain-specific models, underscoring the potential of architectural and training innovations to offset limitations in model scale partially. The key driver is the MeSH-based knowledge hierarchy similarity evaluation integrated within Question Evaluation Agent, which substantially enhances a model’s ability to comprehend and reason about the complex terminologies and conceptual structures of biomedical texts. This demonstrates the effectiveness of our method in adapting smaller, general-purpose models to specialized domains—a critical advantage given the rapid iteration cycles in large model development. In the large-model group, m-KAILIN significantly boosts the performance of LLaMA-3-70B, enabling it to outperform powerful closed-source systems such as GPT-4 with MedPrompt and Med-PaLM-2, despite their larger parameter counts. This suggests that, while increasing model size typically yields better results, targeted preference alignment using knowledge hierarchy evaluation can enable smaller or open-source models to close the gap or even exceed the performance of much larger models, thus offering substantial efficiency and accessibility benefits. In summary, the m-KAILIN framework enables both small and large general-purpose models to achieve state-of-the-art performance on domain-specific biomedical tasks. These experimental results also validate the feasibility of leveraging the m-KAILIN framework to automatically extract high-quality training data from raw textual sources.
4.3 Ablation Studies of each Technical Component
@@ 4.3.1 Ablation Study on Question Generation Agent
4.3.2 Ablation Study on Context Retrieval Agent and Question Evaluation Agent
@@
Figure 9 shows the performance across different configurations. Removing MeSH (w/o MeSH) leads to a substantial decrease in accuracy, especially in the reasoning-required setting. This suggests that biomedical hierarchical knowledge—provided by MeSH—is crucial for the evaluation agent to accurately judge the relevance and quality of context-question pairs. Without this structured guidance, the general LLM struggles to evaluate biomedical content effectively, resulting in lower overall performance. In the w/o Embedding setting, where the retrieval agent uses general-purpose BERT embeddings instead of a fine-tuned, domain-specific model, performance also drops. This indicates that domain-adapted embeddings significantly enhance the retrieval of relevant biomedical contexts. General embeddings may miss domain-specific nuances, leading to suboptimal context selection and thereby affecting downstream QA accuracy. Notably, when both MeSH and domain-adapted embedding are removed (w/o Both), there is a pronounced performance degradation (64.8% in reasoning-required, 43% in question-only), confirming that both hierarchical knowledge in evaluation and domain adaptation in retrieval play complementary, essential roles. Overall, these results demonstrate that integrating domain-specific components—both in evaluation and retrieval—is vital for constructing a high-quality biomedical corpus and maximizing LLM downstream performance.
[…]
@@ 4.4 Hyperparameter Experiment
4.5 Case Study of Generated Corpus
From the specific cases illustrated in Figure 11, it is evident that the model can recognize ”activity” in relation to papain’s function as an enzyme, demonstrating a general understanding facilitated by domain cues. However, without comprehensive and structured biomedical knowledge, the model still exhibits limitations in generating integrative questions that require a deeper or more holistic perspective. Performance differences under the question-only inference setting further highlight these shortcomings. In particular, when the embedding model is not fine-tuned on biomedical data (i.e., using a general-purpose embedding), the model struggles to capture question-specific nuances, resulting in significantly reduced performance when relevant context is unavailable. Our case analysis also reveals the pivotal role of the MeSH-based knowledge hierarchy. When MeSH guidance is present, the model is able to correctly interpret papain as a ”proteolytic enzyme” within the broader biomedical context. However, the absence of fine-grained, domain-specific knowledge—especially when leveraging only general LLMs or embeddings—leads to missed opportunities in question generation, such as failing to inquire about detailed aspects like enzyme activity or dimerization. These findings underscore the necessity of integrating both structured biomedical ontologies (e.g., MeSH) and domain-adapted retrieval mechanisms in our multi-agent framework. Only through this combination can the model attain both the breadth and depth required for high-quality, knowledge-rich biomedical question generation and evaluation.
2020202030303030404040405050505060606060707070708080808090909090100100100100747474747676767678787878808080808282828284848484Scale (×$\times 10^{3}$)Accuracy (%)m-KAILIN-LLaMA-3 (8B)m-KAILIN-LLaMA-3 (70B)
Figure 12: Evaluation (accuracy(%)) of Llama-3-8B and Llama-3-70B on PubMedQA using datasets of varying scales, from 20,000 (20×$\times 10^{3}$) to 100,000 (100×$\times 10^{3}$).
[…]
Figure 13: Evaluations (accuracy (%)) for different chronological subsets of problems in PubMedQA PQA-L on our models compared to other representative models.
[…]
4.6 Scaling Law of Dataset Distillation
We conducted experiments using datasets of varying scales for continuous pre-training, followed by supervised fine-tuning under consistent experimental settings, to evaluate the impact of different data volumes on model performance. In this study, we selected the Llama-3-8B and Llama-3-70B models as base models, to verify the impact of varying dataset scales on models of different sizes. As shown in Figure 12, we observed that the larger LLaMA-3-70B model demonstrates overall better performance compared to the smaller LLaMA-3-8B model. We attributed this advantage to its larger parameter scale, which provides superior feature capture and generalization capabilities. We also observed that on datasets of varying sizes, both models showed a trend of performance improvement as the dataset size increased. The underlying driver is that larger datasets distilled by m-KAILIN are better able to represent a broader range of biomedical knowledge due to a greater number of documents being comprehensively analyzed. This phenomenon indicates that the m-KAILIN framework indeed produces higher-quality datasets when distilling larger-scale datasets. Moreover, we found that while smaller models may not outperform larger models, they seem to benefit more from progressively larger distilled datasets. We attribute this phenomenon to the degree of alignment between task complexity and model scale. This phenomenon indicates that larger models, due to potential knowledge redundancy, may gain less from incremental datasets on a specific task.
4.7 Robustness toward Different Question Groups
4.7.1 Robustness Under Chronological Question
As shown in Figure 13, our agentic framework, especially when combined with MeSH-based evaluation, enables models to maintain consistently high accuracy across documents published at different time periods. Compared to strong biomedical baselines such as MMedLM, models distilled with m-KAILIN demonstrate significantly greater robustness to temporal shifts. This improvement stems from two key innovations: (1) the multi-agent setup, which incorporates diverse perspectives during corpus construction, and (2) the MeSH-guided evaluation agent, which provides structured biomedical context, allowing the model to better understand and adapt to evolving domain knowledge over time. Notably, larger models exhibit even greater temporal robustness, likely due to their increased capacity to absorb and generalize from temporally diverse data curated by our framework.
4.7.2 Robustness Under Subdisciplinary Question
We conduct evaluation on subdisciplinary slices based on major MeSH terms (Figure 14). The results show that m-KAILIN-trained models outperform their peers across various biomedical topics and demographics. This robustness stems from our MeSH-based knowledge-driven design, which leverages MeSH hierarchies during both retrieval and evaluation. This design ensures comprehensive coverage and nuanced understanding across biomedical subfields. The evaluation agent’s use of MeSH facilitates more relevant and context-aware question-answer generation, systematically exposing the model to the breadth of the biomedical literature’s structure. Finally, the overall trend affirms the effectiveness of our approach in producing models well-adapted to the complex, hierarchical landscape of biomedical knowledge.